# Xpikeformer: Hybrid Analog-Digital Hardware Acceleration for Spiking Transformers

### v1

@@ Abstract
This paper introduces Xpikeformer, a hybrid analog-digital hardware architecture designed to accelerate spiking neural network (SNN)-based transformer models. By combining the energy efficiency and temporal dynamics of SNNs with the powerful sequence modeling capabilities of transformers, Xpikeformer leverages mixed analog-digital computing techniques to enhance performance and energy efficiency. The architecture integrates analog in-memory computing (AIMC) for feedforward and fully connected layers, and a stochastic spiking attention (SSA) engine for efficient attention mechanisms. We detail the design, implementation, and evaluation of Xpikeformer, demonstrating significant improvements in energy consumption and computational efficiency. Through an image classification task and a wireless communication symbol detection task, we show that Xpikeformer can achieve software-comparable inference accuracy. Energy evaluations reveal that Xpikeformer achieves up to a 17.817.817.817.819.2×19.2$\times$19.2 × reduction in energy consumption compared to state-of-the-art digital ANN transformers and up to a 5.95.95.95.96.8×6.8$\times$6.8 × reduction compared to fully digital SNN transformers. Xpikeformer also achieves a 12.0×12.0$\times$12.0 × speedup compared to the GPU implementation of spiking transformers.
@@ I Introduction
The transformer architecture has revolutionized various fields within artificial intelligence (AI), becoming the cornerstone of many state-of-the-art (SOTA) models used in natural language processing (NLP) [1], computer vision (CV) [2], and wireless communications [3, 4, 5], but it also imposes a significant burden on computing and memory resources. This is due to their vast number of parameters and the large number of matrix multiplications, particularly during the calculation of the attention mechanism.
@@
The integration of spiking neural networks (SNNs) with transformer architectures represents a promising frontier in the pursuit of energy-efficient and powerful neural computation [6]. SNNs, inspired by the brain’s information processing mechanisms, utilize discrete, temporally-encoded spikes for neuronal signaling [7]. This encoding allows SNNs to achieve higher computational efficiency and lower energy consumption compared to traditional artificial neural networks (ANNs), which rely on continuous, real-valued activations. Spiking transformers, or SNN-based transformers, incorporate spiking dynamics into both feed-forward operations and self-attention mechanisms. They replace traditional multiply-and-accumulate (MAC) operations with accumulate (AC) operations, reducing computational complexity and enhancing efficiency [8, 9, 10, 11, 12, 13, 14, 6, 15, 16].
Although spiking transformers use simplified computation, they still suffer from inefficiencies when deployed on general-purpose computing platforms like CPUs and GPUs. These inefficiencies arise due to two factors: First, spiking transformers experience significant data communication overhead due to frequent memory access for intermediate results, which worsens with longer spike encoding lengths. Second, spiking transformers utilize binary data, while general-purpose computing platforms operate with higher-precision resources, leading to inefficiency in computation and storage [17]. While optimization efforts for spiking transformers have so far focused on algorithm and software levels, the development of dedicated hardware accelerators remains under-researched.
@@
In recent years, analog in-memory computing (AIMC) using non-volatile memory (NVM) or memristor elements has emerged as a promising mixed-signal approach for accelerating both ANNs and SNNs. AIMC enables matrix-vector multiplications (MVMs) to be performed in constant time, $O(1)$, directly within cross-bar arrays of NVM devices. This is achieved by storing the matrix elements as conductances of the NVM devices and by applying input vector values as electrical voltages to the crossbar input lines and reading the resulting electrical currents in the crossbar output lines, leveraging Kirchoff’s laws [18]. Several AIMC-based ANN transformer accelerators have been proposed, significantly improving the efficiency of ANN transformers [19, 20, 21]. However, due to architectural differences between ANN transformers and spiking transformers, these accelerators cannot be directly applied to the latter.
Additionally, writing data to memristor crossbars introduces extra latency and energy consumption, and the memristor devices suffer from limited endurance, i.e., they can only support a limited number of read/write cycles before catastrophic failure[22]. Therefore, AIMC is primarily useful for accelerating the inference process in feedforward or fully connected layers, where weights are written to the crossbar once and reused for forward propagation. In the attention mechanism of spiking transformers, the two matrices involved in the matrix multiplication are time-varying activations, making AIMC unsuitable for accelerating this process as the intermediate result storage can quickly exceed the endurance limit [19]. Consequently, designing a new hardware architecture to accelerate the spike-based attention mechanism is necessary.
@@
In this paper, we introduce Xpikeformer, a hybrid analog-digital hardware solution specifically designed to accelerate spiking transformers. This architecture combines the strengths of two distinct computing engines to optimize performance and energy efficiency. The AIMC engine handles static-weight operations such as feed-forward, embedding, and fully connected layers, by utilizing specialized memory arrays for efficient data storage and processing. For the attention mechanism, Xpikeformer employs the stochastic spiking attention (SSA) engine, which utilizes a simplified computing approach based on logical operations and a streaming dataflow, ensuring fast and efficient processing without the need to store intermediate results.
By integrating these two specialized engines, Xpikeformer offers a robust and efficient solution tailored to the unique computational needs of spiking transformers, effectively addressing the inefficiencies found in general-purpose computing platforms. The key contributions of this work are as follows:
  • We present Xpikeformer, the first hardware architecture for spiking transformers, comprising an AIMC engine and an SSA engine. We also introduce a simulation framework to train and evaluate the accuracy performance of Xpikeformer.
@@
  • We design spiking neuron tiles within the AIMC engine to efficiently handle feed-forward, embedding, and fully connected layers without storing non-binary pre-activations, thereby significantly reducing data transfer overhead.
@@
  • We introduce a software-hardware co-design methodology for spike-based attention architecture, SSA, that builds on stochastic computing and is specifically designed to execute matrix multiplications in the attention mechanism of spiking transformers using lightweight logic gates rather than heavy-weight arithmetic units. We evaluate the proposed SSA tile on an FPGA, demonstrating 18×18$\times$18 × less latency while consuming 17×17$\times$17 × less power compared to traditional ANN attention on a GPU.
  • We evaluate the energy consumption of the two tasks tested, Xpikeformer achieved a 17.817.817.817.819.2×19.2$\times$19.2 × energy reduction compared to the state-of-the-art digital ANN transformer accelerator, a 5.05.05.05.07.1×7.1$\times$7.1 × energy reduction compared to the hybrid implementation of ANN, and a 5.95.95.95.96.8×6.8$\times$6.8 × energy reduction compared to the fully digital implementation of spiking transformers on general-purpose devices. Xpikeformer also achieves a 1.53×1.53$\times$1.53 × speedup than traditional ANN on GPU and a 12.0×12.0$\times$12.0 × speedup compared to the GPU implementation of spiking transformers.
@@
This paper is structured as follows: Section II provides an overview of the preliminaries of the work. Section III discusses related work on hardware accelerators for ANN and SNN transformers. Section IV details the design and implementation of the Xpikeformer architecture. Section V covers the training and inference methodologies. Section VI presents the accuracy evaluation of the proposed system, and Section VII evaluates its efficiency. Finally, Section VIII concludes the paper.
@@ II Preliminaries
Figure 1: (a) Illustration of an artificial neuron (left) and a spiking neuron (right). (b) A simplified block diagram of an SNN transformer.
II-A Spike Coding and Spiking Neurons
One essential difference between SNNs and ANNs is the information encoding method. In ANNs, activations are represented by real values $x$. In contrast, SNNs encode activations in binary temporal sequences $s[t]\in\{0,1\}$ for $t=1,2,\ldots,T$, known as spike trains, where $T$ denotes the spike encoding length. A commonly utilized spike encoding scheme to map a real value $x\in[0,1]$ into a spike train is Bernoulli rate coding, whereby each bit in the encoded sequence independently takes the value ‘1’ with probability $x$ and ‘0’ with probability $1-x$, which is expressed as:
$s[t]\sim\text{Bern}(x)\text{ for }t=1,2,\ldots,T.$ (1)
$V^{t}$ $=\beta V^{t-1}+I^{t},$ (2)
where $\beta\in(0,1]$ represents the leak factor determining the rate of decay of the membrane potential over time, and $I^{t}$ is the input at time step $t$. When $V^{t}$ reaches or exceeds a threshold voltage $V_{\text{thresh}}$, the neuron fires a spike, i.e., $O^{t}=1$, resetting the potential $V^{t}$ to 00. Otherwise, the neuron remains silent, producing $O^{t}=0$, i.e.,
[…]
@@ II-B Stochastic Computing
An important advantage of Bernoulli rate coding is that it supports stochastic computing, a computational paradigm that employs Bernoulli bit streams [23, 24]. Consider two Bernoulli spike trains $s_{1}[t]$ and $s_{2}[t]$ with spiking probabilities $p\left(s_{1}[t]=1\right)=x_{1}$ and $p\left(s_{2}[t]=1\right)=x_{2}$. The multiplication of real-valued $x_{1}$ and $x_{2}$ can be achieved by applying the bit-wise logic AND ($\land$) operation on their stochastic representations $s_{1}[t]$ and $s_{2}[t]$, which is denoted as:
@@ II-C Transformers and Spiking Transformers
II-C1 ANN Transformer
The typical architecture of an ANN transformer model is based on stacks of encoder layers [25, 2], decoder layers [26], or both [1]. Each encoder or decoder layer is composed of a multi-head self-attention (MHSA) mechanism and a feedforward network with residual connections and layer normalization. In models that include both encoder and decoder stacks, the decoder layers also contain an additional multi-head self-attention mechanism that attends to the encoder’s output.
@@ II-C2 Spiking Transformers
Spiking transformers are a novel adaptation of traditional ANN transformers that integrates principles from SNNs [8, 9, 10, 11, 12, 13, 14, 6, 15, 16]. A typical implementation of a spiking transformer is depicted in Fig. 1. In this model, input tokens are initially encoded into temporal spike trains through a spike encoding layer. The subsequent architecture mirrors that of ANN transformers, with the operations of ANN transformers and spiking transformers compared in Table I. The operation of the proposed Xpikeformer is also listed in Table I and will be detailed in Section IV-B.
For a spiking transformer, the feedforward, embedding, and fully connected layers are implemented using spiking neurons in spiking transformers. In the context of MHSA, a notable difference lies in how the token representations $Q$, $K$, and $V$ are generated and processed. In ANN transformers, the representations $Q$, $K$, and $V$ are obtained through linear projections applied to the token embeddings $X$. In contrast, spiking transformers apply LIF neurons after linear projections, resulting in spike-encoded representations $Q^{t}$, $K^{t}$, and $V^{t}$. Compared to the attention score computation $S=\text{softmax}(QK^{\mathsf{T}}/\sqrt{d_{K}})$ in ANN transformers, spiking transformers first perform matrix multiplication between matrices $Q^{t}$ and $K^{t}$ independently for each time step $t$, and then apply a LIF neuron to the product, which is expressed as $S^{t}=\text{LIF}(Q^{t}{K^{t}}^{\mathsf{T}})$. A notable difference in spiking transformer MHSA is the absence of softmax operations. Traditional ANN transformers use the softmax function to normalize the attention scores, ensuring they sum to one. However, in SNNs, the use of softmax is impractical due to the binary nature of spike trains and the need for temporal and sparse processing. Instead, the spiking neurons inherently handle normalization through their spiking activity [10]. The attention score $S^{t}$ is then multiplied with $V^{t}$ for every $t$ independently. Then a LIF neuron is applied to obtain the final attention results.
@@ II-C3 Efficiency Challenges
Spiking transformers have demonstrated comparable accuracy to their ANN counterparts in fields such as NLP [13], computer vision [12], and wireless communication [27]. These models can provide up to an order of magnitude gain in computational energy efficiency by simplifying the complex floating-point (FP) matrix multiplications of conventional implementations into addition and integer multiplications. However, implementing spike-based models on general-purpose computing platforms, such as CPUs and GPUs, generally leads to significant energy inefficiencies due to the following reasons:
  1. 1.
    Data Communication Overhead: Spiking transformers do not mitigate the data communication overhead caused by frequent memory access required for storing and reading intermediate results within and between feedforward and MHSA layers. Although intermediate results are mostly one-bit data, this overhead can be exacerbated when the temporal dimension (spike encoding length) exceeds the bit width used in ANNs. Additionally, SNNs generate non-binary pre-activations at every time step, leading to transmission overheads that are $T$ times larger than those in ANNs with the same quantization level, where $T$ is the spike encoding length.
  2. 2.
    Precision Mismatch: SNNs generate and compute with a large amount of binary data, but general-purpose computing platforms typically use high-precision resources for computation and storage. For instance, CPUs typically operate with 32-bit or 64-bit floating-point (FP32/64) precision. While some GPUs can compute using as low as 8-bit integers (INT8), this is still significantly over-provisioned for handling binary spiking signals. The development of spiking attention accelerators on FPGA devices could potentially address this issue, but this area remains under-researched [28].
@@ II-D Analog In-Memory Computing (AIMC)
AIMC enables fast and energy-efficient MVMs using synaptic arrays (SAs). An SA primarily comprises a memristor crossbar. The crossbars consist of memristor devices, which are NVM elements positioned at the intersections of rows and columns that can be programmed to different conductance levels, representing the matrix elements. This configuration allows the crossbar to store a matrix with significantly greater area efficiency compared to traditional digital buffers. To perform an MVM with a crossbar, the real-valued vector is encoded as input voltages applied to the rows of the crossbar. Ohm’s Law governs the current through each memristor, and Kirchhoff’s current law (KCL) ensures the correct summation of these currents at the columns, corresponding to the dot-product result. Beyond the crossbar, an SA also includes readout units, which typically comprise successive approximation-register (SAR) analog-to-digital converters (ADCs) coupled with multilevel current-mode sense amplifiers. These are coupled with accumulation circuits, such as adders and registers, as well as various peripheral components, including decoders, multiplexers, switch matrices, and buffers.
@@
A typical implementation of an SA, introduced by [29], is illustrated in Fig. 2. In this setup, every two devices in the crossbar form a differential pair cell, allowing the representation of positive and negative matrix values. The input spike-encoded signals are sent from the bit lines (BLs), and the results are read from the source lines (SLs). The switching on and off of the memristor is controlled by the word line (WL) and SL through a transistor. During the weighted sum operation, all cells are activated when all WLs are turned on. Input vector voltages are applied to the BLs, and the resulting weighted sum currents are read out through the SLs in parallel. These currents are then digitized by multiple readout units concurrently. To optimize the balance between area, power, and latency, the readout units are shared among a predefined number of crossbar columns and time-multiplexed under the control of a multiplexer (MUX).
In a spiking transformer, input neuron activations are represented by spike trains, which consist of multiple cycles of input voltage signals to the BLs. Because these input signals represent 1-bit binary data at each time step, there is no need to use analog voltages to represent them, thus eliminating the need for digital-to-analog converters (DACs) during inference. However, DACs would need to be provisioned in the hardware to permit the initial programming of the memristor conductances to the required analog levels corresponding to the matrix elements.

### v2

@@ Abstract
The integration of neuromorphic computing and transformers through spiking neural networks (SNNs) offers a promising path to energy-efficient sequence modeling, with the potential to overcome the energy-intensive nature of the artificial neural network (ANN)-based transformers. However, the algorithmic efficiency of SNN-based transformers cannot be fully exploited on GPUs due to architectural incompatibility. This paper introduces Xpikeformer, a hybrid analog-digital hardware architecture designed to accelerate SNN-based transformer models. The architecture integrates analog in-memory computing (AIMC) for feedforward and fully connected layers, and a stochastic spiking attention (SSA) engine for efficient attention mechanisms. We detail the design, implementation, and evaluation of Xpikeformer, demonstrating significant improvements in energy consumption and computational efficiency. Through image classification tasks and wireless communication symbol detection tasks, we show that Xpikeformer can achieve inference accuracy comparable to the GPU implementation of ANN-based transformers. Evaluations reveal that Xpikeformer achieves 13×13$\times$13 × reduction in energy consumption at approximately the same throughput as the state-of-the-art (SOTA) digital accelerator for ANN-based transformers. Additionally, Xpikeformer achieves up to 1.9×1.9$\times$1.9 × energy reduction compared to the optimal digital ASIC projection of SOTA SNN-based transformers.
@@ I Introduction
The transformer architecture has revolutionized various fields within artificial intelligence, becoming the backbone of many state-of-the-art (SOTA) models used in natural language processing (NLP) [1], computer vision (CV) [2], and wireless communications [3, 4, 5], but it also imposes a significant burden on computing and memory resources. This is due to their vast number of parameters and the large amount of matrix-vector and matrix-matrix multiplications required by dense feed-forward layers as well as the attention mechanism.
@@
The integration of neuromorphic computing with transformer architectures, enabled by spiking neural networks (SNNs), represents a promising frontier for achieving energy-efficient and biologically-plastic sequence modeling [6, 7, 8, 9]. As illustrated in Fig. 1, neuromorphic computing uses sparse, temporally-encoded spikes for neuronal signaling [10]. This spike-based signaling allows spiking neurons to aggregate activations using efficient accumulate (AC) operations, in contrast to artificial neurons in traditional artificial neural networks (ANNs), which process continuous-valued activations through multiply-and-accumulate (MAC) operations. SNN-based transformers, or spiking transformers, as shown in Fig. 1, adopts spiking neurons in feed-forward operations while incorporating spiking-based self-attention mechanisms, achieving higher computing efficiency in sequence modeling tasks compared to ANN-based transformers [11, 12, 13, 14, 15, 16, 17, 6, 18, 19].
Figure 1: (a) Comparison of an artificial neuron (left) and a spiking neuron (right). (b) A simplified block diagram of a spiking transformer.
Although spiking transformers are algorithmically efficient due to their simplified computations, they face significant inefficiencies on general-purpose computing platforms like GPUs. These inefficiencies stem from a fundamental mismatch between the event-driven, sparse nature of spiking algorithms and GPU architectures, which are optimized for dense, high-precision workloads. First, spiking transformers operate on binary data, while GPUs are designed for high-precision arithmetic (e.g., FP16 or FP32), leading to poor resource utilization and suboptimal energy efficiency. Second, the temporal dimension of spiking transformers introduces substantial data communication overhead, as frequent memory access for intermediate results significantly increases latency and energy consumption. While optimization efforts for spiking transformers have primarily focused on algorithmic and software improvements, the lack of dedicated hardware accelerators remains a critical gap in the field [20].
@@
In recent years, analog in-memory computing (AIMC) using non-volatile memory (NVM) devices has emerged as a promising mixed-signal approach for accelerating both ANNs and SNNs [21]. AIMC enables matrix-vector multiplications (MVMs) to be performed in constant time, $O(1)$, directly within crossbar arrays of NVM devices. This is achieved by storing the matrix elements as conductances of the NVM devices and by applying input vector values as electrical voltages to the crossbar input lines and reading the resulting electrical currents in the crossbar output lines, leveraging Kirchhoff’s laws. Several AIMC-based ANN transformer accelerators have been proposed, significantly improving the efficiency of ANN transformers [22, 23, 24]. However, due to architectural differences between ANN-based transformers and spiking transformers, these accelerators cannot be directly applied to the latter.
Additionally, AIMC is primarily effective for accelerating inference in feedforward and fully connected layers, where weights are programmed into the crossbar once and reused during forward propagation. However, AIMC is less suited for attention mechanisms, which involve matrix multiplications between two dynamic matrices, requiring frequent data writes to NVM crossbars. This introduces latency, increases energy overhead, and is limited by NVM endurance [25, 22]. As a result, digital acceleration is better suited for attention. Although these matrix multiplications can be simplified to masked additions in spiking domain owing to the binary nature of the multiplicands [13], executing them on GPUs—designed for high-precision arithmetic—remains resource-inefficient.
@@
In this paper, we introduce Xpikeformer, the first hardware accelerator specifically designed for spiking transformers. The architecture integrates two specialized computing frameworks to optimize performance and energy efficiency. For static-weight operations, including feed-forward, embedding, and fully connected layers, Xpikeformer employs AIMC for efficient in-memory computation and weight storage. For the attention mechanism, it utilizes stochastic computing with lightweight logical operations and a streaming dataflow, allowing rapid processing without intermediate result storage. Additionally, we introduce a simulation framework to train Xpikeformer and evaluate its accuracy performance. The key contributions of this work are as follows:
@@
  • We design an AIMC engine to efficiently handle feed-forward, embedding, and fully connected layers in spiking transformers without storing non-binary pre-activations, significantly reducing data transfer overhead.
@@
  • We propose a software-hardware co-design methodology for a spike-based attention architecture, termed stochastic spiking attention (SSA), which leverages stochastic computing to execute matrix multiplications in the attention mechanism of spiking transformers using lightweight logic gates instead of heavy arithmetic units.
  • Xpikeformer achieves a 13×13$\times$13 × energy reduction at approximately the same throughput as the SOTA digital accelerator for ANN-based transformer. As the first hardware accelerator design for spiking transformers, Xpikformer also showcases 1.81.81.81.81.9×1.9$\times$1.9 × energy reduction compared to the ideal digital ASIC projection of the SOTA spiking transformer software design. Compared to GPU implementations, Xpikeformer achieves a 2.18×2.18$\times$2.18 × speedup over traditional ANN-based transformers and a 6.85×6.85$\times$6.85 × speedup over spiking transformers.
@@
This paper is structured as follows: Section II provides an overview of the preliminaries. Section III reviews related work on hardware accelerators for ANN-based and spiking transformers. Section IV details the design and implementation of the Xpikeformer architecture. Section V discusses the training and inference methodologies. Section VI presents the accuracy evaluation of the proposed system, and Section VII assesses its efficiency. Finally, Section VIII concludes the paper.
@@ II Preliminaries
II-A Spike Encoding and Spiking Neurons
One essential difference between SNNs and ANNs is the information encoding method [26]. In ANNs, activations are represented by real values $x$. In contrast, SNNs encode activations in binary temporal sequences $s[t]\in\{0,1\}$ for $t=1,2,\ldots,T$, known as spike trains, where $T$ denotes the spike encoding length. A commonly utilized spike encoding scheme to map a real value $x\in[0,1]$ into a spike train is Bernoulli rate coding, whereby each bit in the encoded sequence independently takes the value ‘1’ with probability $x$ and ‘0’ with probability $1-x$, which is expressed as:
$s[t]\sim\text{Bern}(x)\text{ for }t=1,2,\ldots,T.$ (1)
$V^{t}$ $=\beta V^{t-1}+I^{t},$ (2)
where $\beta\in(0,1]$ represents the leak factor determining the rate of decay of the membrane potential over time, and $I^{t}$ is the input at time step $t$. When $V^{t}$ reaches or exceeds a threshold voltage $V_{\text{thresh}}$, the neuron fires a spike, i.e., $O^{t}=1$, and resets the potential $V^{t}$ to 00. Otherwise, the neuron remains silent, producing $O^{t}=0$, i.e.,
[…]
@@ II-B Stochastic Computing
An important advantage of Bernoulli rate coding is that it supports stochastic computing, a computational paradigm that employs Bernoulli bit streams [27]. Consider two Bernoulli spike trains $s_{1}[t]$ and $s_{2}[t]$ with spiking probabilities $p\left(s_{1}[t]=1\right)=x_{1}$ and $p\left(s_{2}[t]=1\right)=x_{2}$. The multiplication of real-valued $x_{1}$ and $x_{2}$ can be achieved by applying the bit-wise logic AND ($\land$) operation on their stochastic representations $s_{1}[t]$ and $s_{2}[t]$, which is denoted as:
@@ II-C Transformers and Spiking Transformers
II-C1 ANN-based Transformer
The typical architecture of an ANN-based transformer model is based on stacks of encoder layers [28, 2], decoder layers [29], or both [1]. Each encoder or decoder layer is composed of a multi-head self-attention (MHSA) mechanism and a feedforward network with residual connections and layer normalization. In models that include both encoder and decoder stacks, the decoder layers also contain an additional multi-head self-attention mechanism that attends to the encoder’s output.
@@ II-C2 Spiking Transformers
Spiking transformers are a novel adaptation of traditional ANN-based transformers that integrates principles from SNNs [11, 12, 13, 14, 15, 16, 17, 6, 18, 19]. The typical architecture of a spiking transformer is shown in Fig. 1. In this architecture, input tokens are first encoded into temporal spike trains through a spike encoding layer. The subsequent architecture resembles that of ANN transformers, with the operations of ANN-based transformers and spiking transformers compared in Table I. The operation of the proposed Xpikeformer is also listed in Table I and will be detailed in Section IV-B.
In a spiking transformer, the feedforward, embedding, and fully connected layers are implemented using spiking neurons in spiking transformers. A key distinction in MHSA lies in how the token representations $Q$, $K$, and $V$ are generated and processed. In ANN-based transformers, the representations $Q$, $K$, and $V$ are obtained through linear projections applied to the token embeddings $X$. In contrast, spiking transformers apply LIF neurons after linear projections, resulting in spike-encoded representations $Q^{t}$, $K^{t}$, and $V^{t}$. Compared to the attention score computation $S=\text{softmax}(QK^{\mathsf{T}}/\sqrt{d_{K}})$ in ANN-based transformers, spiking transformers first perform matrix multiplication between matrices $Q^{t}$ and $K^{t}$ independently for each time step $t$, then apply a LIF neuron to the product, expressed as $S^{t}=\text{LIF}(Q^{t}{K^{t}}^{\mathsf{T}})$. Another notable difference in spiking transformer MHSA is the absence of softmax. ANN-based transformers use the softmax function to normalize the attention scores, ensuring they sum to one. However, in spiking transformers, softmax is impractical due to the binary nature of attention scores and the need for temporal and sparse processing. Instead, spiking neurons regulate the activation through their firing dynamics [13]. Finally, the attention score $S^{t}$ is multiplied with $V^{t}$ at each time step $t$, followed by a LIF neuron to produce the final attention output.
@@ II-C3 Efficiency Challenges
Spiking transformers achieve accuracy comparable to their ANN counterparts in NLP [16], CV [15], and wireless communication [30]. Estimated on neuromorphic hardware assumptions, their energy efficiency suggests up to an order-of-magnitude reduction in computational energy by replacing matrix multiplications with addition and integer multiplication. However, deploying spike-based models on general-purpose platforms like CPUs and GPUs results in significant energy inefficiencies due to:
  • Data Communication Overhead: Frequent memory access for intermediate results in feedforward and MHSA layers introduces overhead, especially when the spike encoding length $T$ exceeds GPU bit widths. Non-binary pre-activations generated at each time step further amplify this overhead, increasing transmission requirements by a factor of $T$ compared to equally quantized ANNs.
  • Precision Mismatch: SNNs operate on binary data, whereas CPUs and GPUs are optimized for high-precision computation (e.g., FP32/64), leading to inefficiencies. While some GPUs support INT8, even this precision exceeds the needs of binary spiking signals. FPGA-based spiking attention accelerators may address this issue, but remain underexplored [31].
@@ II-D Analog In-Memory Computing (AIMC)
AIMC enables fast and energy-efficient MVMs using synaptic arrays (SAs). An SA primarily comprises an NVM crossbar. The crossbars consist of NVM devices, which are NVM elements positioned at row-column intersections, where each element can be programmed to different conductance levels, representing matrix values. This structure allows the crossbar to store matrices with significantly greater area efficiency than traditional digital buffers. To perform an MVM, the input vector is encoded as voltages applied to the crossbar rows. Ohm’s Law governs the current flow through each NVM device, while Kirchhoff’s Current Law ensures the correct summation of currents at the columns, yielding the dot-product result. Beyond the crossbar, an SA includes readout units, typically comprising successive approximation-register (SAR) analog-to-digital converters (ADCs) coupled with multilevel current-mode sense amplifiers. These are integrated with accumulation circuits (e.g., adders and registers) and various peripheral components, including decoders, multiplexers, switch matrices, and buffers.
A typical implementation of an SA, introduced by [32], is illustrated in Fig. 2. In this setup, every two devices in the crossbar form a differential pair cell, enabling the representation of positive and negative matrix values. The input spike-encoded signals are applied through the bit lines (BLs), while the results are read from the source lines (SLs). The word lines (WLs) and SLs control the switching of NVM devices via transistors. During the weighted sum operation, all cells are activated when all WLs are turned on. Input vector voltages are applied to the BLs, and the resulting weighted sum currents are read out in parallel through the SLs. These currents are then digitized by multiple readout units concurrently. To optimize the balance between area, power, and latency, the readout units are shared among a predefined number of crossbar columns and time-multiplexed under the control of a multiplexer (MUX).
In a spiking transformer, input neuron activations are represented by spike trains, which consist of multiple cycles of input voltage signals to the BLs. Since these input signals encode 1-bit binary data at each time step, analog voltage representation is unnecessary, eliminating the need for digital-to-analog converters (DACs) during inference. However, DACs are still required in hardware for the initial programming of NVM conductances to the appropriate analog levels corresponding to the matrix elements.